# Gaussian Building Mesh (GBM): Extract a Building’s 3D Mesh with Google Earth and Gaussian Splatting

### v2

@@ Abstract
Recently released open-source pre-trained foundational image segmentation and object detection models (SAM2+GroundingDINO) allow for geometrically consistent segmentation of objects of interest in multi-view 2D images. Users can use text-based or click-based prompts to segment objects of interest without requiring labeled training datasets. Gaussian Splatting allows for the learning of the 3D representation of a scene’s geometry and radiance based on 2D images. Combining Google Earth Studio, SAM2+GroundingDINO, 2D Gaussian Splatting, and our improvements in mask refinement based on morphological operations and contour simplification, we created a pipeline to extract the 3D mesh of any building based on its name, address, or geographic coordinates.
Index Terms:
Remote Sensing, Gaussian Splatting, SAM, 3D Mesh, Building
I Introduction
@@
Leveraging Google Earth Studio [1], and inspired by GS2Mesh[2], we propose a 3D building mesh extraction pipeline capable of extracting the 3D mesh of a building given its proper name, address, postal code, or geographical coordinates. Our contributions are as follows:
  • We combine Google Earth Studio, the Segment Anything Model, and a Gaussian Splatting mesh extraction model into a 3D building mesh extraction pipeline. This allows for the generation of the 3D mesh of a building from its name, address, postal code, or latitude/longitude coordinates.
  • We leverage Segment Anything Model-2[3] with GroundingDINO[4] for 3D consistent building masking with either text or click-based masking input modalities, improving on GS2Mesh-based mask extraction. We add mask re-prompting and refinement as contributions.
  • We created a mask refinement method based on morphological operations and the Ramer-Douglas-Peucker algorithm.
  • We use an improved version of 2DGS as the Gaussian Splatting model and generate the 3D colored mesh of a building of interest using masked Truncated Signed Distance Function (TSDF) Fusion/Integration. During the meshing process, we make many minor but effective adjustments, such as filtering and smoothing the depth maps and adjusting the hyperparameters during both Gaussian Splatting training and TSDF meshing.
Google Earth StudioSAM2 Mask ExtractionMask Refinement2D Gaussian SplattingTSDF FusionRemote sensing imagesImages and masksImages and refined masksDepth and color mapsSynthesized 2D images3D colored mesh
@@
GS2Mesh [2] combines out-of-the-box object segmentation and 3DGS to generate 3D meshes of objects regardless of the background. To our knowledge, it is the only method similar to our own and the source of our inspiration. Compared to GS2Mesh, we use a state-of-the-art (not-yet-published) fork of 2DGS [5] instead of the original 3DGS method for 3D representation learning. This allows us to obtain accurate depth maps without using deep learning-based stereo depth reconstruction, rendering GS2Mesh’s DLNR stereo depth reconstruction module obsolete, which we remove. This also allows for another significant change: because we no longer need a background as context for DLNR, we move the automated building masking step to before Gaussian splatting training, i.e., we train Gaussian splatting on masked images as opposed to full images. This increases the training speed by a factor of $\sim$5, depending on the portion of the region masked as background. We also add mask re-prompting mask refinements based on morphological operations and contour simplification, which correct for SAM2 masking errors, as well as depth smoothing, which helps with TSDF meshing. These improvements greatly increased the mesh extraction accuracy and efficiency. Our pipeline is shown in Figure 1 and detailed in Section IV.
II Related Works
II-A 3D Reconstruction from 2D Images
Classical 3D reconstruction from 2D images is a well-studied area. Photogrammetry is well-understood in remote sensing and computer vision and is widely used for academic, industrial, and commercial purposes. For this purpose, multiple software suites and code libraries [6, 7, 8, 9, 10, 11] are commercially available or open-sourced, with some specifically designed for remote sensing applications. Many of these not only allow for the extraction of a sparse point cloud from 2D images but also generate a 3D mesh. These 3D meshes are often more desirable than point clouds, as they are more photorealistic and allow for use in 3D modeling and simulations. COLMAP [6], in particular, is worth highlighting. It is a photogrammetry and Structure-from-Motion (SfM) library widely used as a preprocessing step for modern novel view synthesis and learning-based 3D rendering methods and is included in most Gaussian Splatting models’ pipelines, including our own.
II-B Modern Novel-View Synthesis and Neural/Learning-based Rendering
We categorize two broad families of recently developed methods in this section: NeRF and Gaussian Splatting, with a focus on papers that apply these techniques to remote sensing and building model extraction. Given the extensive body of work in this field, it is more comprehensively captured in survey papers [12, 13] than in a few paragraphs of literature review.
@@ II-B Modern Novel-View Synthesis and Neural/Learning-based Rendering
Neural Radiance Field (NeRF) [14], introduced in 2020, has garnered significant attention in this field. In NeRF-based models, the 3D representation consists of a 3D radiance field (directionally dependent color field) and a 3D density field, both represented as Multi-Layer Perceptrons (MLPs). Differentiable volume rendering is employed to generate 2D images by sampling and integrating local 3D color/radiance and density. These radiance and density fields are trained from scratch using a photometric loss function, meaning NeRF models learn a 3D representation of a scene from 2D images. Significant advancements have been made in this area, including [15, 16, 17], with models like [18, 19, 20] focusing on improving 3D geometry extraction, and others like [21, 22, 23, 24, 25] applying NeRF techniques to remote sensing or urban scene capture.
@@
3D Gaussian Splatting (3DGS) [26], first proposed in 2023, and subsequent Gaussian Splatting models have largely surpassed NeRF-based approaches over the past year. 3DGS uses a large number of 3D Gaussian distributions (also known as primitives in computer graphics) as a learned 3D representation. In addition to standard 3D Gaussian distribution function parameters, each Gaussian primitive also has directionally dependent color and opacity $\alpha$, which are all trainable parameters. The method uses a differentiable tile-based rasterizer, projecting the 3D Gaussians onto the to-be-rasterized image, and $\alpha$-blending the projected Gaussians. This allows the 3D Gaussians representing the scene’s radiance and geometry to be learned from scratch using a photometric loss function. Compared to NeRF, Gaussian Splatting models are generally much faster to train, have higher view-synthesis quality, but require more memory. Methods such as [27, 28, 29] improved 3D extraction, and methods such as [30, 31, 32, 33, 34] applied Gaussian Splatting to remote sensing or urban scene capture.
@@
However, we note that these methods all focus on scene-wide capture and cannot extract the 3D mesh of individual objects without further processing. Only GS2Mesh [2], being an out-of-the-box pipeline leveraging Segment Anything Model object masking, allows for the extraction of a 3D mesh from user prompting without further training or processing. Unfortunately, according to our preliminary testing, GS2Mesh often fails in remote sensing building extraction scenarios due to issues in the mask generation module and the pretrained deep learning-based stereo depth map module, which motivated our research.
III Background
III-A 3D Gaussian Splatting
@@ III Background
3D Gaussian Splatting (3DGS) [26] is a view-synthesis technique that enables the learning of a 3D scene’s geometry and lighting from multi-view 2D images, which can then be used to rasterize the scene from novel viewpoints. The process begins with an initial point cloud, often generated using COLMAP [6] Structure-from-Motion (SfM). For each point in this cloud, 3DGS initializes a Gaussian primitive that encodes learnable parameters such as mean, covariance, opacity, and local lighting in the red, green, and blue channels, represented as spherical harmonic coefficients.
@@
Figure 2: 2D Gaussian Splatting uses surfaced aligned 2D Gaussian primitives embedded in 3D to represent the 3D scene. 2D Gaussian is represented by it’s 3D position $\mathbf{p_{k}}$, it’s scale $s_{u},s_{v}$, and it’s orientation $\mathbf{t_{u},t_{v}}$. Figure 3 in [27]
@@
2D Gaussian Splatting (2DGS) [27] enhances the standard 3D Gaussian Splatting (3DGS), improving the reproduction of 3D surface geometry. While GS2Mesh slightly outperforms 2DGS on the DTU benchmark, the two methods are fundamentally different. GS2Mesh is primarily a mesh extraction pipeline that utilizes vanilla 3DGS during the 3D reconstruction phase. In contrast, 2DGS is a significant improvement on 3DGS itself, altering the nature of the Gaussian splats. As a result, 2DGS can be used to replace 3DGS in many pipelines, as we have done.
@@
The key innovation in 2D Gaussian Splatting is the representation of the scene using 2D-oriented planar Gaussians instead of 3D Gaussians. Like standard 3DGS, 2DGS employs Gaussian primitives that store spherical harmonic coefficients for each color channel, local transparency $\alpha$, and 3D location $\mathbf{p}_{k}$. However, unlike 3D Gaussian primitives, 2D Gaussian primitives have two scalar values to represent variance ($s_{u},s_{v}$) and two tangent vectors ($\mathbf{t}_{u},\mathbf{t}_{v}$) whose cross product results in the normal vector that defines orientation (see figure 2). Depth maps can be accurately rendered using the projected depth value. For more details, we refer readers to the original paper [27].
III-C Segment Anything Model (version 2) and Grounding DINO
@@
Segment Anything Model (SAM) [35] is an out-of-the-box image segmentation model pretrained on a massive billion-image dataset. It is capable of segmenting most objects given point-based or bounding box priors, without requiring further training from the user. Users can provide point-click or bounding box prompts to identify the object(s) of interest, and SAM will return segmentation masks and associated scores. However, a key limitation of SAM is that, when applied to a scene viewed from multiple images, the individual masks generated are not necessarily consistent with each other. This inconsistency limits SAM’s effectiveness in segmenting video data, which requires temporal consistency, and multi-view data, which requires 3D consistency.
@@
SAM2 [3], released in August 2024, addresses this issue by introducing consistent video segmentation that maintains 3D and temporal consistency through the use of memory attention. Since we aim to extract the mesh of an individual building rather than an entire neighborhood, segmentation masks are crucial.
@@
GroundingDINO [4] is a pretrained open-set object detector capable of extracting object bounding boxes in images from natural language prompts without requiring additional training from the user. GroundingDINO can be combined with SAM/SAM2 to enable text-based object segmentation. The process involves first generating a bounding box from the text description and then using the bounding box to prompt SAM/SAM2. This combination of GroundingDINO and SAM/SAM2 is known as the Grounded-SAM pipeline [36], which is available as an open-source library.
III-D GS2Mesh
@@
GS2Mesh [2] is a Gaussian Splatting-based 3D reconstruction pipeline, outperforming concurrent and competing methods such as SuGAR [28], 2DGS [27], and GOF [29] on the DTU dataset [37] benchmark.
@@
  • The entire ensemble of depth images are integrated into a mesh using the Truncated Signed Distance Function fusion (TSDF) algorithm [38] with the Marching-Cubes algorithm [39].
@@
From these stereo pairs, DLNR [40], pretrained on the Middlebury dataset [41], is used for depth extraction from stereo images. To improve the quality of the reconstructions, multiple masks are applied to the stereo model’s output to filter out regions visible to only one camera and to discard depth estimates outside the valid range.
@@
Following depth extraction, a standard Open3D [42] implementation of the TSDF algorithm initializes and populates a voxel grid with the scene geometry. The voxel representation of the scene is populated with the signed distance to the nearest scene surface, integrated from the depth images generated by DLNR. The marching cubes algorithm then assigns each cube’s vertices in the voxel grid to be inside or outside the nearest surface based on the previously calculated TSDF values. Based on the 8-vertex configuration, a local surface is meshed for each cube, which is repeated across the entire voxel grid. This process generates a mesh from the voxel representation.
III-E Evaluation Metrics
@@
For 2D view synthesis visual quality assessment, we use the commonly accepted Peak Signal-to-Noise Ratio (PSNR) [43], 2D Structural Similarity Index Measure (SSIM) [44], and Learned Perceptual Image Patch Similarity (LPIPS) [45]. These are full-reference metrics that compare an assessed image with a ground truth image. PSNR and SSIM are higher when the assessed image and the ground truth are similar. SSIM achieves a maximum value of 1 when the two images are identical. LPIPS, on the other hand, is lower when the two images are similar, with a minimum of 0 when the two images are identical.
@@
For 3D mesh quality assessment, we use the 3D Structural Similarity Index Measure (3D-SSIM) [46], comparing a 360-degree rendering video of the mesh with a 360-degree ground truth video created by segmenting the building from its background in the Google Earth training images. We note that there are other full-reference 3D geometry and visual quality metrics that compare 3D models to other 3D models. However, we lack ground truth 3D models for the buildings we meshed and only have access to ground truth 2D images. This is the main reason for using the video-based 3D-SSIM for mesh quality comparison. We provide both the average 3D-SSIM across the entire video and the minimum 3D-SSIM across video frames.
III-F Google Earth Studio

### v3

# Investigation, Writing - Review & Editing

@@ Abstract
The rapid convergence of computer vision, and digital technologies is redefining how buildings are captured, modeled, and managed. In computer vision, Recently released open-source pre-trained foundational image segmentation and object detection models allow for geometrically consistent segmentation of objects of interest in multi-view 2D images. Text-based or click-based prompts can be used to segment objects of interest without requiring labeled training datasets, allowing for both user-prompted or automated segmentation. Simultaneously, Gaussian Splatting allows for learning a 3D representation of a scene’s geometry and radiance based on 2D images. Combining Google Earth Studio, SAM2+GroundingDINO, 2D Gaussian Splatting, and our improvements in mask refinement based on morphological operations and contour simplification, we created a pipeline to extract the 3D mesh of any building based on its name, address, or geographic coordinates. Our pipeline offers a fast and user-accessible for rapid 3D modeling of built environments and structures, enabling downstream applications.
keywords:
Remote Sensing; Gaussian Splatting; SAM; 3D Mesh; 3D Building Model
1 Introduction
@@
Leveraging Google Earth Studio (Alphabet Inc., 2015-2024), and inspired by GS2Mesh(Wolf et al., 2025), we propose a 3D building mesh extraction pipeline capable of extracting the 3D mesh of a building given its proper name, address, postal code, or geographical coordinates. When combined with off-the-shelf registration methods, our pipeline enables downstream tasks such as building information models, construction verification, and automated safety assessment.
Our contributions are as follows:
  • We create a novel meshing pipeline that allows for the extraction of a 3D mesh of a building from its name, address, postal code, or latitude/longitude coordinates without using on-site LiDAR, or camera data.
  • We leverage and improve Segment Anything Model-2 (Ravi et al., 2024) and GroundingDINO (Liu et al., 2025) for geometrically consistent building masking. We add mask refinements based on morphological operations and the Ramer-Douglas-Peucker algorithm combined with mask re-prompting, which can aid in other use cases of SAM2 and GroundingDINO.
  • We improved and modified an unpublished implementation of 2DGS (Huang et al., 2024) to generate 3D colored building meshes via masked TSDF integration, with refinements in depth map filtering, smoothing, and hyperparameter tuning during both training and meshing.
GS2Mesh (Wolf et al., 2025) combines out-of-the-box object segmentation and 3DGS to generate 3D meshes of objects regardless of the background. To our knowledge, it is the only method similar to our own that performs object-based 3D mesh extraction from text prompts or click prompts. This method was the main source of our inspiration. There are many other Gaussian Splatting-based 3D geometry and mesh extraction methods. However, they focus on reconstructing the entire scene as opposed to being able to extract the mesh of designated objects based on user-based text input, clicks, or system-level automation.
Google Earth StudioOne of {address, place name, postal code, geographic coordinates}SAM2 Mask ExtractionMask Refinement2D Gaussian SplattingTSDF FusionRemote sensing imagesImages and masksImages and refined masksDepth and color mapsSynthesized images3D colored mesh
@@
2 Related Works
2.1 3D Reconstruction from 2D Images
Classical 3D reconstruction from 2D images is a well-studied area. Photogrammetry is well-understood in remote sensing and computer vision and is widely used for academic, industrial, and commercial purposes. For this purpose, multiple software suites and code libraries (Schonberger and Frahm, 2016; Snavely, 2008; Adorjan, 2016; Reality, 2024; Solutions, 2024; Esri, 2024) are commercially available or open-sourced, with some specifically designed for remote sensing applications. Many of these not only allow for the extraction of a sparse point cloud from 2D images but also generate a 3D mesh. These 3D meshes are often more desirable than point clouds, as they are more photorealistic and allow for use in 3D modeling and simulations. COLMAP (Schonberger and Frahm, 2016), in particular, is worth highlighting. It is a photogrammetry and Structure-from-Motion (SfM) library widely used as a preprocessing step for modern novel view synthesis and learning-based 3D rendering methods and is included in most Gaussian Splatting models’ pipelines, including our own.
2.2 Modern Novel-View Synthesis and Neural/Learning-based Rendering
We categorize two broad families of recently developed methods in this section: NeRF and Gaussian Splatting, with a focus on papers that apply these techniques to remote sensing and building model extraction. Given the extensive body of work in this field, it is more comprehensively captured in survey papers (Gao et al., 2022; Chen and Wang, 2024) than in a few paragraphs of literature review.
@@ 2.2 Modern Novel-View Synthesis and Neural/Learning-based Rendering
Neural Radiance Field (NeRF) (Mildenhall et al., 2021), introduced in 2020, has garnered significant attention in this field. In NeRF-based models, the 3D representation consists of a 3D radiance field (directionally dependent color field) and a 3D density field, both represented as Multi-Layer Perceptrons (MLPs). Differentiable volume rendering is employed to generate 2D images by sampling and integrating local 3D color/radiance and density. These radiance and density fields are trained from scratch using a photometric loss function, meaning NeRF models learn a 3D representation of a scene from 2D images. Significant advancements have been made in this area, including (Barron et al., 2021, 2022; Müller et al., 2022), with models like (Wang et al., 2021; Oechsle et al., 2021; Wang et al., 2023) focusing on improving 3D geometry extraction, and others like (Xiangli et al., 2022; Rematas et al., 2022; Tancik et al., 2022; Derksen and Izzo, 2021; Marí et al., 2022) applying NeRF techniques to remote sensing or urban scene capture.
@@
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), first proposed in 2023, and subsequent Gaussian Splatting models have largely surpassed NeRF-based approaches over the past year. 3DGS uses a large number of 3D Gaussian distributions (also known as primitives in computer graphics) as a learned 3D representation. In addition to standard 3D Gaussian distribution function parameters, each Gaussian primitive also has directionally dependent color and opacity $\alpha$, which are all trainable parameters. The method uses a differentiable tile-based rasterizer, projecting the 3D Gaussians onto the to-be-rasterized image, and $\alpha$-blending the projected Gaussians. This allows the 3D Gaussians representing the scene’s radiance and geometry to be learned from scratch using a photometric loss function. Compared to NeRF, Gaussian Splatting models are generally much faster to train, have higher view-synthesis quality, but require more memory. Methods such as (Huang et al., 2024; Guédon and Lepetit, 2024; Yu et al., 2024) improved 3D extraction, and methods such as (Zhou et al., 2024; Li et al., 2025; Wang et al., 2024; Chen et al., 2024; Gao et al., 2024a) applied Gaussian Splatting to remote sensing or urban scene capture.
@@
However, we note that these methods all focus on scene-wide capture and cannot extract the 3D mesh of individual objects without further processing. Only GS2Mesh (Wolf et al., 2025), being an out-of-the-box pipeline leveraging Segment Anything Model object masking, allows for the extraction of a 3D mesh from user prompting without further training or processing. Unfortunately, according to our preliminary testing, GS2Mesh often fails in remote sensing building extraction scenarios due to issues in the mask generation module and the pretrained deep learning-based stereo depth map module, which motivated our research.
3 Background
3.1 3D Gaussian Splatting
@@ 3 Background
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) is a view-synthesis technique that enables the learning of a 3D scene’s geometry and lighting from multi-view 2D images, which can then be used to rasterize the scene from novel viewpoints. The process begins with an initial point cloud, often generated using COLMAP (Schonberger and Frahm, 2016) Structure-from-Motion (SfM). For each point in this cloud, 3DGS initializes a Gaussian primitive that encodes learnable parameters such as mean, covariance, opacity, and local lighting in the red, green, and blue channels, represented as spherical harmonic coefficients.
@@
Figure 2: 2D Gaussian Splatting uses surfaced aligned 2D Gaussian primitives embedded in 3D to represent the 3D scene. 2D Gaussian is represented by it’s 3D position $\mathbf{p_{k}}$, it’s scale $s_{u},s_{v}$, and it’s orientation $\mathbf{t_{u},t_{v}}$. (Huang et al., 2024)
@@
2D Gaussian Splatting (2DGS) (Huang et al., 2024) enhances the standard 3D Gaussian Splatting (3DGS), improving the reproduction of 3D surface geometry. While GS2Mesh slightly outperforms 2DGS on the DTU benchmark, the two methods are fundamentally different. GS2Mesh is primarily a mesh extraction pipeline that utilizes vanilla 3DGS during the 3D reconstruction phase. In contrast, 2DGS is a significant improvement on 3DGS itself, altering the nature of the Gaussian splats. As a result, 2DGS can be used to replace 3DGS in many pipelines, as we have done.
@@
The key innovation in 2D Gaussian Splatting is the representation of the scene using 2D-oriented planar Gaussians instead of 3D Gaussians. Like standard 3DGS, 2DGS employs Gaussian primitives that store spherical harmonic coefficients for each color channel, local transparency $\alpha$, and 3D location $\mathbf{p}_{k}$. However, unlike 3D Gaussian primitives, 2D Gaussian primitives have two scalar values to represent variance ($s_{u},s_{v}$) and two tangent vectors ($\mathbf{t}_{u},\mathbf{t}_{v}$) whose cross product results in the normal vector that defines orientation (see figure 2). Depth maps can be accurately rendered using the projected depth value. For more details, we refer readers to the original paper (Huang et al., 2024).
3.3 Segment Anything Model (version 2) and Grounding DINO
@@
Segment Anything Model (SAM) (Kirillov et al., 2023) is an out-of-the-box image segmentation model pretrained on a massive billion-image dataset. It is capable of segmenting most objects given point-based or bounding box priors, without requiring further training from the user. Users can provide point-click or bounding box prompts to identify the object(s) of interest, and SAM will return segmentation masks and associated scores. However, a key limitation of SAM is that, when applied to a scene viewed from multiple images, the individual masks generated are not necessarily consistent with each other. This inconsistency limits SAM’s effectiveness in segmenting video data, which requires temporal consistency, and multi-view data, which requires 3D consistency.
@@
SAM2 (Ravi et al., 2024), released in August 2024, addresses this issue by introducing consistent video segmentation that maintains 3D and temporal consistency through the use of memory attention. Since we aim to extract the mesh of an individual building rather than an entire neighborhood, segmentation masks are crucial.
@@
GroundingDINO (Liu et al., 2025) is a pretrained open-set object detector capable of extracting object bounding boxes in images from natural language prompts without requiring additional training from the user. GroundingDINO can be combined with SAM/SAM2 to enable text-based object segmentation. The process involves first generating a bounding box from the text description and then using the bounding box to prompt SAM/SAM2. This combination of GroundingDINO and SAM/SAM2 is known as the Grounded-SAM pipeline (Ren et al., 2024), which is available as an open-source library.
3.4 GS2Mesh
@@
GS2Mesh (Wolf et al., 2025) is a Gaussian Splatting-based 3D reconstruction pipeline, outperforming concurrent and competing methods such as SuGAR (Guédon and Lepetit, 2024), 2DGS (Huang et al., 2024), and GOF (Yu et al., 2024) on the DTU dataset (Sølund et al., 2016) benchmark.
@@
  • The entire ensemble of depth images are integrated into a mesh using the Truncated Signed Distance Function fusion (TSDF) algorithm (Curless and Levoy, 1996) with the Marching-Cubes algorithm (Lorensen and Cline, 1998).
@@
From these stereo pairs, DLNR (Zhao et al., 2023), pretrained on the Middlebury dataset (Scharstein and Szeliski, 2002), is used for depth extraction from stereo images. To improve the quality of the reconstructions, multiple masks are applied to the stereo model’s output to filter out regions visible to only one camera and to discard depth estimates outside the valid range.
@@
Following depth extraction, a standard Open3D (Zhou et al., 2018) implementation of the TSDF algorithm initializes and populates a voxel grid with the scene geometry. The voxel representation of the scene is populated with the signed distance to the nearest scene surface, integrated from the depth images generated by DLNR. The marching cubes algorithm then assigns each cube’s vertices in the voxel grid to be inside or outside the nearest surface based on the previously calculated TSDF values. Based on the 8-vertex configuration, a local surface is meshed for each cube, which is repeated across the entire voxel grid. This process generates a mesh from the voxel representation.
3.5 Evaluation Metrics
@@
For 2D view synthesis visual quality assessment, we use the commonly accepted Peak Signal-to-Noise Ratio (PSNR) (Netravali, 2013), 2D Structural Similarity Index Measure (SSIM) (Wang et al., 2004), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018). These are full-reference metrics that compare an assessed image with a ground truth image. PSNR and SSIM are higher when the assessed image and the ground truth are similar. SSIM achieves a maximum value of 1 when the two images are identical. LPIPS, on the other hand, is lower when the two images are similar, with a minimum of 0 when the two images are identical.
@@
For 3D mesh quality assessment, we use the 3D Structural Similarity Index Measure (3D-SSIM) (Zeng and Wang, 2012), comparing a 360-degree rendering video of the mesh with a 360-degree ground truth video created by segmenting the building from its background in the Google Earth training images. We note that there are other full-reference 3D geometry and visual quality metrics that compare 3D models to other 3D models. However, we lack ground truth 3D models for the buildings we meshed and only have access to ground truth 2D images. This is the main reason for using the video-based 3D-SSIM for mesh quality comparison. We provide both the average 3D-SSIM across the entire video and the minimum 3D-SSIM across video frames.
3.6 Google Earth Studio